During FY 2019, Dorchester Company plans to sell Widgets for $13 a unit. Current variable costs are $7 a unit and fixed costs are expected to total of $136,000. Use this information to determine the number of units of Widgets for Dorchester to breakeven. (Round to the nearest whole number)
Answer #1

Contribution margin=Sales-Variable costs

=(13-7)=$6 per unit

Hence breakeven sales=Fixed cost/Contribution margin

=(136000/6)

=22,667 units(Approx).